OpenStudy (mlgpro):

Determine which data set would have the largest mean. 2, 8, 12, 16 1, 9, 12, 18 1, 8, 12, 28 8, 9, 12, 28

OpenStudy (kropot72):

The mean is the sum of all the values in the data set, divided by the number of values in the data set. So the mean of the first data set is: \[\frac{2+8+12+16}{4}=\frac{38}{4}=9.5\] There are four values in each data set, therefore the set with the largest sum of its values will have the largest mean.

OpenStudy (landshark5):

do u know how to calculate the mean

OpenStudy (mlgpro):

no @kropot72 did not help me

OpenStudy (landshark5):

first you add all the numbers. all the data. For example: 2,8,12,16. You add those numbers to get 38. With me so far?

OpenStudy (mlgpro):

yup

OpenStudy (landshark5):

Then you divide by the amount of the data. In this case you have 4 pieces of data since the're 4 numbers. So 38/4. That gives you 9.5. The mean of the first set of data is 9.5

OpenStudy (landshark5):

any problems?

OpenStudy (landshark5):

basically you add all the numbers and divide by the amount of numbers.

OpenStudy (landshark5):

want another example?

OpenStudy (mlgpro):

hold on i have another question for u

OpenStudy (landshark5):

I got a better example if you want

OpenStudy (landshark5):

ok

OpenStudy (landshark5):

do you get the problem though?

OpenStudy (mlgpro):

yeah i jus cant find the 3rd set`s mean

OpenStudy (landshark5):

1+8+12+28=49

OpenStudy (mlgpro):

/4=?

OpenStudy (landshark5):

there are 4 sets of data so u divide by 4

OpenStudy (mlgpro):

12.25?

OpenStudy (landshark5):

49/4=12.25

OpenStudy (landshark5):

yes

OpenStudy (mlgpro):

then the 4th 47/4=i do not know

OpenStudy (landshark5):

that is 11.75

OpenStudy (mlgpro):

so it is the third

OpenStudy (landshark5):

yes

OpenStudy (landshark5):

all you gotta do is add the numbers. and them divide by the amount of numbers.

OpenStudy (mlgpro):

do you know how to do measures of central tendancy

OpenStudy (landshark5):

well let me see the problems first

OpenStudy (landshark5):

might or might not, Its been 3 years since 6th grade.

OpenStudy (landshark5):

are u in 6th?

OpenStudy (mlgpro):

1.Which measure of central tendency best describes this data set? 52, 58, 53, 57, 91 Mean Median Mode Range 2.The weight, in pounds, of 6 small dogs is: 16, 20, 17, 16, 17, 16. A somewhat overweight dog joins them. The new dog weighs 28 pounds. Which measure of central tendency is not affected by placing this new dog into the group? Mean Median Mode None of the above 3.The number of unpopped kernels in 7 bags of popcorn is: 13, 17, 8, 11, 10, 13, 15. If each value is doubled, what would happen to the mean amount of unpopped kernels for these 7 bags? It would not change. It would double. It would increase by two. It would be cut in half.

OpenStudy (anonymous):

The last choice D would have the largest mean. Mean = Average.

OpenStudy (landshark5):

yeah i for got the last one forgot to calculate that

OpenStudy (yanasidlinskiy):

Hint: Median = Middle. Set the ordered numbers from least to greatest and find out which one is in the middle:D

OpenStudy (landshark5):

my mistake

OpenStudy (landshark5):

mlg: you have to see if the median changes of you add those numbers.

OpenStudy (landshark5):

u there?

OpenStudy (mlgpro):

yea

OpenStudy (mlgpro):

@landshark5 yeah my dog needed to get put in the bath sorry